Business and finance. The cost in dollars of manufacturing w wing nuts is given by the expression 0.07w + 13.3. Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?
Answer by tutorcecilia(2152)

C(w)=0.07w+13.3 [substitute (w=375) and solve]
C(375)=0.07(375)+13.3
C(375)=39.55
.
C(w)=0.07w+13.3 [substitute (w=1) and solve]
C(1)=0.07(1)+13.3
C(1)=13.79